OpenStudy (piercetheveil47):

List some example for me below ? :) 1.)Write your own real-world story problem that requires adding or subtracting mixed numbers with unlike denominators. 2.)Solve your problem from Question 1. Show all of your work for full credit.

OpenStudy (solomonzelman):

restaurant owner. I bought 8 cakes for 8 parties, I cut 2 cakes into 3 parts because it meant to be served for 3 people, and similarly I cut 3 more cakes into 7 parts, and other 3 cakes into 8 parts. Then someone came and ate a piece from each cake. How many piece are there left ?

OpenStudy (solomonzelman):

So I cut the cakes cake 1 3 / 3 cake 2 3 / 3 cake 3 7 / 7 cake 4 7 / 7 cake 5 7 / 7 cake 6 8 / 8 cake 7 8 / 8 cake 8 8 / 8

OpenStudy (solomonzelman):

So I cut the cakes then someone eats/takes 1 peace from each cake 1 3 / 3 - 1/3 = 2/3 of the cake cake 2 3 / 3 - 1/3 = 2/3 of the cake cake 3 7 / 7 - 1/7 = 6/7 of the cake cake 4 7 / 7 - 1/7 = 6/7 of the cake cake 5 7 / 7 - 1/7 = 6/7 of the cake cake 6 8 / 8 - 1/8 = 7/8 of the cake cake 7 8 / 8 - 1/8 = 7/8 of the cake cake 8 8 / 8 - 1/8 = 7/8 of the cake

OpenStudy (solomonzelman):

So I have left 3 × (7 / 8) from the cakes I cut for 3 different 8-people parties 3 × (6 /7) from the cakes I cut for 3 different 7-people parties 2 × (6 /7) from the cakes I cut for 2 different 3-people parties (By the way, my question should be how many cakes are there left after all)

OpenStudy (solomonzelman):

So, so far I am going to get \(\Large\color{blue}{ 3\frac{7}{8} + 3\frac{6}{7}+2\frac{2}{3} }\)

OpenStudy (solomonzelman):

Then solve the question. \(\Large\color{blue}{ 3\frac{7}{8} + 3\frac{6}{7}+2\frac{2}{3} }\) Can you add it?

OpenStudy (solomonzelman):

I have "Sabbath" (Or Shabat ) Coming soon, and I am an Orthodox Jew, so i am going offline soon. Good Luck !

OpenStudy (piercetheveil47):

yes thank you so much this was a great example @SolomonZelman & have a good rest of your evening!

OpenStudy (solomonzelman):

I would sign, to brag "creative teenager"... just kidding. You (will) see that it is not so hard to make something like this up.... Have a good day:)

OpenStudy (piercetheveil47):

you too thank you! @SolomonZelman